# For Better or For Worse? Learning Minimum Variance Features With Label Augmentation

### v1

@@ 5 Conclusion
In summary, we have shown that for linearly separable data distributions in which there exist high variance features and much lower variance features, linear models trained with the data augmentation schemes of label smoothing and Mixup fail to learn the low variance features in both practice and theory, while standard training with weight decay succeeds. The main practical takeaway from this is that label smoothing and Mixup do not necessarily always improve the robustness of models, and their failure to do so may be indicative of certain identifying features in the training data (i.e. adversarially introduced patches). A natural follow-up to the results we have presented is an extension to the case of non-linear models and non-linearly-separable data, which the results of Section 4.3 indicate should be possible.
@@ Acknowledgements
Appendix A Full Proofs of Main Results
Here we provide full proofs of all of the results contained in Section 3 of the main paper.
A.1 Proof of Theorem 3.1 and Corollary 3.2
See 3.1
@@ Appendix A Full Proofs of Main Results
First, we claim that $w_{\mathcal{L}}^{*}$ is in the direction of $v_{1}^{*}$, which is the normal vector to the max-margin hyperplane separating $v_{1}^{*}$ and $v_{-1}^{*}$ (since $v_{1}^{*}=-v_{-1}^{*}$ due to the fact that the data distribution has mean zero and $\pi_{Y}(1)=\pi_{Y}(-1)$). Indeed, for any other choice of $w\in\mathbb{R}^{d}$ and any $(x,y)\in\mathrm{supp}(\pi)$, we have:
$y\left\langle w_{\mathcal{L}},x_{\mathcal{L}}\right\rangle=y\left\langle w_{\mathcal{L}},v^{*}_{y}\right\rangle=\left\langle w_{\mathcal{L}},v^{*}_{1}\right\rangle$ (A.1)
As a result, if $w_{\mathcal{L}}$ is not in the direction of $v_{1}^{*}$, we can remove the component orthogonal to $v_{1}^{*}$ (thereby decreasing the norm of the solution) while keeping the inner product with $x_{\mathcal{L}}$ the same. Now let $u_{1}^{*}=v_{1}^{*}/\norm{v_{1}^{*}}$ and $u_{2}^{*}\in\mathbb{R}^{\absolutevalue{\mathcal{H}}}$ similarly denote the unit normal vector to the max-margin hyperplane separating $A_{1}$ and $A_{-1}$. Recall that $\left\langle u_{2}^{*},x_{\mathcal{H}}\right\rangle>\left\langle u_{1}^{*},x_{\mathcal{L}}\right\rangle$ by assumption three in Definition 2.1.
Let us identify $u_{1}^{*}$ and $u_{2}^{*}$ with the corresponding vectors in $\mathbb{R}^{d}$ whose $-\mathcal{L}$ and $-\mathcal{H}$ components are zero respectively, so that $u_{1}^{*}$ and $u_{2}^{*}$ are orthonormal in $\mathbb{R}^{d}$. Then we can decompose the unique minimizer of the loss $w^{*}$ as $\alpha u_{1}^{*}+\beta u_{2}^{*}+\kappa u_{3}^{*}$, where $u_{3}^{*}$ is a vector orthonormal to $u_{1}^{*}$ and $u_{2}^{*}$.
Now suppose that $\alpha>\beta$. Then we claim that replacing $\alpha$ and $\beta$ with $\gamma=\sqrt{(\alpha^{2}+\beta^{2})/2}$ yields a solution with lower loss than $w^{*}$.
To see this, first observe that $2\gamma^{2}=\alpha^{2}+\beta^{2}$, so that the norm of the modified solution is the same as $w^{*}$. This implies that the weight decay penalty term in $\ell_{\lambda}$ is unchanged.
Furthermore, we have that $\gamma\in(\beta,\alpha)$, and that $\gamma-\beta>\alpha-\gamma$. This follows from the fact that $\gamma\geq(\alpha+\beta)/2$ by construction. This, combined with the fact that $y\left\langle u_{2}^{*},x_{\mathcal{H}}\right\rangle>y\left\langle u_{1}^{*},x_{\mathcal{L}}\right\rangle$, then implies for all $(x,y)\in\mathrm{supp}(\pi)$ that:
$y\left\langle\gamma u_{1}^{*}+\gamma u_{2}^{*},x\right\rangle>y\left\langle\alpha u_{1}^{*}+\beta u_{2}^{*},x\right\rangle$ (A.2)
Which contradicts the minimality of $w^{*}$. Therefore, we must have $\alpha\leq\beta$, which gives the desired result. ∎
See 3.2
@@
The strategy of the proof will be to construct a stationary point for the label smoothed loss that is correlated only with the dimensions in $\mathcal{L}$, and then argue using Jensen’s inequality that any solution that is correlated with the dimensions in $\mathcal{H}$ will have non-zero variance and thereby worse loss than our constructed stationary point. We first compute the gradient $\gradient_{w}\ell_{\mathrm{LS},\alpha}(w)$ and Hessian $\gradient_{w}^{2}\ell_{\mathrm{LS},\alpha}(w)$:
$\gradient_{w}\ell_{\mathrm{LS},\alpha}(w)$ $=\mathbb{E}_{(x,y)\sim\pi}\left[(g_{w}(yx)-1+\alpha/2)yx\right]$ (A.3)
$\gradient_{w}^{2}\ell_{\mathrm{LS},\alpha}(w)$ $=\mathbb{E}_{x\sim\pi_{X}}\left[g_{w}(yx)(1-g_{w}(yx))xx^{\top}\right]$ (A.4)
Equations (A.3) and (A.4) follow from interchanging the gradient operator with the expectations, which is justified in this case via the dominated convergence theorem because $g_{w}(x)$ is continuously differentiable (with bounded derivative) in each component of $w$, and $x$ has bounded support. Clearly the Hessian is positive semidefinite, so $\ell_{\mathrm{LS},\alpha}(w)$ is convex.
It is also straightforward to construct a stationary point $w^{*}$ of $\ell_{\mathrm{LS},\alpha}(w)$ by choosing $w^{*}_{\mathcal{L}}=\beta v_{1}^{*}$ for appropriate $\beta\in\mathbb{R}$ and $w^{*}_{\mathcal{H}}=0$, since this allows for $g_{w^{*}}(v_{1}^{*})=1-\alpha/2$ and $g_{w^{*}}(-v_{1}^{*})=g_{w^{*}}(v_{-1}^{*})=\alpha/2$ (here $v_{1}^{*}=-v_{-1}^{*}$ since $\pi_{X}$ has mean zero). Now we claim that any $w^{\prime}$ with $\norm{w^{\prime}_{\mathcal{H}}}>0$ has strictly greater loss than $w^{*}$.
To show this, we first observe that $\ell_{\mathrm{LS},\alpha}$ is also convex as a function of $w^{\top}x$ (this is easily check from Equations (A.3) and (A.4)). Additionally, we can restrict our attention to $w^{\prime}$ satisfying $\norm{w^{\prime}}\leq M$ for some $M$, since $\ell_{\mathrm{LS},\alpha}(w^{\prime})\to\infty$ as $\norm{w^{\prime}}\to\infty$ for $\alpha>0$. Therefore, we can consider $w^{\top}x$ to be integrable and apply the conditional form of Jensen’s inequality to $\ell_{\mathrm{LS},\alpha}(w)$:
$(\displaystyle\ell_{\mathrm{LS},\alpha}(w)\geq\sum_{y\in\{-1,1\}}-\pi_{Y}(y)\bigg{(}$$\bigg{(}1-\frac{\alpha}{2}\bigg{)}\log\sigma\big{(}\mathbb{E}[yw^{\top}x\mid Y=y]\big{)}$
$+\frac{\alpha}{2}\log\sigma\big{(}-\mathbb{E}[yw^{\top}x\mid Y=y]\big{)}\bigg{)}$ (A.5)
The RHS of Equation (A.5) is minimized by our choice of $w^{*}$, and in fact we have equality between $\ell_{\mathrm{LS},\alpha}(w^{*})$ and the RHS. Furthermore, since we are dealing with a non-affine function of $w^{\top}x$, we can only achieve equality if $w^{\top}x$ is almost surely constant conditioned on $Y=y$. However, by assumption two in Definition 2.1, the conditional variance of $(w_{\mathcal{H}}^{\prime})^{\top}x_{\mathcal{H}}$ is non-zero, and therefore $\ell_{\mathrm{LS},\alpha}(w^{\prime})>\ell_{\mathrm{LS},\alpha}(w^{*})$. ∎
A.3 Proof of Theorem 3.4
See 3.4
@@
We largely follow the same structure as the proof of Theorem 3.3. As before, we compute $\gradient_{w}\ell_{\mathrm{MIX},\mathcal{D}_{\lambda}}(w)$ and $\gradient_{w}^{2}\ell_{\mathrm{MIX},\mathcal{D}_{\lambda}}(w)$ (once again by applying dominated convergence):
$\gradient_{w}\ell_{\mathrm{MIX},\mathcal{D}_{\lambda}}(w)$ $=\mathbb{E}_{(x_{1},y_{1}),(x_{2},y_{2})\sim\pi}\left[\mathbb{E}_{\lambda\sim\mathcal{D}_{\lambda}}\left[\left(\lambda(g_{w}(y_{1}z_{\lambda})-1)y_{1}+(1-\lambda)(g_{w}(y_{2}z_{\lambda})-1)y_{2}\right)z_{\lambda}\right]\right]$ (A.6)
$\gradient_{w}^{2}\ell_{\mathrm{MIX},\mathcal{D}_{\lambda}}(w)$ $=\mathbb{E}_{(x_{1},y_{1}),(x_{2},y_{2})\sim\pi}\bigg{[}\mathbb{E}_{\lambda\sim\mathcal{D}_{\lambda}}\bigg{[}\bigg{(}\lambda(1-g_{w}(y_{1}z_{\lambda}))g_{w}(y_{1}z_{\lambda})$
$\quad\quad\quad\quad\quad\quad+(1-\lambda)(1-g_{w}(y_{2}z_{\lambda}))g_{w}(y_{2}z_{\lambda})\bigg{)}z_{\lambda}z_{\lambda}^{\top}\bigg{]}\bigg{]}$ (A.7)
It follows from Equation (A.7) that the Hessian is positive semidefinite (since $\lambda\in[0,1]$), so $\ell_{\mathrm{MIX},\mathcal{D}_{\lambda}}(w)$ is also convex.
Now it also follows from the above calculations that, after conditioning on $\lambda,y_{1},\text{ and }y_{2}$, the function $h_{\lambda}$ from Equation (2.8) (redefined below to indicate that $y_{1},y_{2}$ are also fixed) is convex in $w^{\top}z_{\lambda}$:
$h_{\lambda,y_{1},y_{2}}(w,z_{\lambda})=-\left(\lambda\log g_{w}(y_{1}z_{\lambda})+(1-\lambda)\log g_{w}(y_{2}z_{\lambda})\right)$ (A.8)
For convenience, we will use $z_{\lambda}\mid y_{1},y_{2}$ to indicate the conditional distribution of $\lambda x_{1}+(1-\lambda)x_{2}$ given $\lambda,y_{1},y_{2}$ as in Equation (A.8). For the moment we suppose that $w^{\top}z_{\lambda}$ is integrable. Then by Fubini’s theorem, conditioning, and Jensen’s inequality we obtain:
$\ell_{\mathrm{MIX},\mathcal{D}_{\lambda}}(w)$ $=\mathbb{E}_{\lambda\sim\mathcal{D}_{\lambda}}\left[\mathbb{E}_{y_{1},y_{2}\sim\pi_{Y}}\left[\mathbb{E}_{z_{\lambda}\mid y_{1},y_{2}}\left[h_{\lambda,y_{1},y_{2}}(w,z_{\lambda})\right]\right]\right]$
$\geq\mathbb{E}_{\lambda\sim\mathcal{D}_{\lambda}}\left[\mathbb{E}_{y_{1},y_{2}\sim\pi_{Y}}\left[-\left(\lambda\log\sigma(y_{1}\mathbb{E}[w^{\top}z_{\lambda}])+(1-\lambda)\log\sigma(y_{2}\mathbb{E}[w^{\top}z_{\lambda}])\right)\right]\right]$ (A.9)
Equality is only obtained in Equation (A.9) if and only if $w^{\top}z_{\lambda}\mid y_{1},y_{2}$ is almost surely constant, which only occurs if $\norm{w_{\mathcal{H}}}=0$ as before. It remains to show in this case that we can obtain a stationary point with $\norm{w_{\mathcal{H}}}=0$, as then that will imply that any choice of $w$ with $\norm{w_{\mathcal{H}}}>0$ will have strictly worse loss.
To do so, let us consider $w$ with $\norm{w_{\mathcal{H}}}=0$ and $\left\langle w,v_{1}^{*}\right\rangle=u=-\left\langle w,v_{-1}^{*}\right\rangle$. Let $\mathbb{P}_{\lambda}$ denote the probability measure associated with $\mathcal{D}_{\lambda}$. We then define the following two functions, which will correspond to the terms in $\gradient_{w}\ell_{\mathrm{MIX},\mathcal{D}_{\lambda}}(w)$ that result from mixing $y_{1}=y_{2}$ and $y_{1}\neq y_{2}$.
$f_{1}(u)$ $=\sigma(u)-1$ (A.10)
$f_{2}(u,\lambda)$ $=\sigma((2\lambda-1)u)-\lambda$ (A.11)
Now, recalling that $\mu_{y}=\mathbb{E}[X\mid Y=y]$ and that $\mu_{1}=-\mu_{-1}$ since $\pi_{Y}(1)=\pi_{Y}(-1)$ and $\mathbb{E}_{x\sim\pi_{X}}[x]=0$, we get:
$\gradient_{w}\ell_{\mathrm{MIX},\mathcal{D}_{\lambda}}(w)$ $=\mathbb{E}_{\lambda\sim\mathcal{D}_{\lambda}}\left[\mathbb{E}_{y_{1},y_{2}\sim\pi_{Y}}\left[\mathbb{E}_{z_{\lambda}\mid y_{1},y_{2}}\left[\left(\lambda(g_{w}(y_{1}z_{\lambda})-1)y_{1}+(1-\lambda)(g_{w}(y_{2}z_{\lambda})-1)y_{2}\right)z_{\lambda}\right]\right]\right]$
$=\mathbb{E}_{\lambda\sim\mathcal{D}_{\lambda}}\left[\frac{1}{2}\mathbb{E}_{z_{\lambda}\mid y_{1}=1,y_{2}=1}\left[f_{1}(u)z_{\lambda}\right]+\frac{1}{2}\mathbb{E}_{z_{\lambda}\mid y_{1}=1,y_{2}=-1}\left[f_{2}(u,\lambda)z_{\lambda}\right]\right]$
@@
Where in the penultimate step we used the fact that the terms that result from conditioning with $y_{1}=y_{2}$ are the same after substituting $\mu_{-1}=-\mu_{1}$, and likewise for the mixed term. Now we consider the limiting behavior of Equation (A.12) as $u\to\pm\infty$. We first focus on the integral term in Equation (A.12), which we can analyze via dominated convergence:
@@
$\quad+\int_{0.5}^{1}\big{(}\sigma((2\lambda-1)u)-\lambda\big{)}(2\lambda-1)\ d\mathbb{P}_{\lambda}(\lambda)\bigg{)}$
[…]
$\geq 0$ (A.13)
And similarly, the limit as $u\to-\infty$ is bounded above by 0. Thus, from Equation (A.12) we get that:
$\lim_{u\to\infty}\gradient_{w}\ell_{\mathrm{MIX},\mathcal{D}_{\lambda}}(w)$ $=\beta_{1}\mu_{1}\quad\text{for some}\quad\beta_{1}\geq 0$ (A.14)
$\lim_{u\to-\infty}\gradient_{w}\ell_{\mathrm{MIX},\mathcal{D}_{\lambda}}(w)$ $=\beta_{2}\mu_{1}\quad\text{for some}\quad\beta_{2}\leq 0$ (A.15)
Since Equation (A.12) is continuous in $u$, it then follows that there exists $u$ such that $\gradient_{w}\ell_{\mathrm{MIX},\mathcal{D}_{\lambda}}(w)=0$, which implies that we can make our choice of $w$ a stationary point. However, there is one subtlety here - the fact that any other $w^{\prime}$ with $\norm{w^{\prime}_{\mathcal{H}}}>0$ has worse loss hinges on the use of Jensen’s inequality above, which requires that we constrain ourselves to the case where $w^{\top}z_{\lambda}$ is integrable.
[…]
For this case, if $\norm{w^{\prime}_{\mathcal{H}}}>0$, then $\ell_{\mathrm{MIX},\mathcal{D}_{\lambda}}(w^{\prime})\to\infty$ as $\norm{w^{\prime}}\to\infty$, since we incur infinite loss on the mixed terms $y_{1}\neq y_{2}$. As a result, we can conclude that every minimizer of $\ell_{\mathrm{MIX},\mathcal{D}_{\lambda}}$ (for $\mathcal{D}_{\lambda}$ satisfying the conditions of the theorem) has zero norm in the high variance dimensions, including those minimizers that occur as the norm of the weights tend to infinity. ∎
Appendix B Low Variance Experiments
(a) Weight Decay
(b) Label Smoothing
(c) Mixup
Figure 6: Final weight norm of logistic regression model in the high variance dimensions ($\norm{w_{\mathcal{H}}}$) for various hyperparameter settings on the synthetic data distribution of Section 4.1 modified so that the dimensions in $\mathcal{L}$ have non-zero variance.
(a) Weight Decay
(b) Label Smoothing
(c) Mixup
Figure 7: Logistic regression final test errors for various hyperparameter settings on the binary classification versions of CIFAR-10 and CIFAR-100 from Section 4.2, but with the training data modified to have non-zero variance in the identifying dimension.
Figures 6 and 7 correspond to non-zero variance versions of the experiments of Sections 4.1 and 4.2. In Figure 6, we consider synthetic data as per Definition 4.1 except for that $x_{i}=y(\gamma+\xi)$ where $\xi\sim\mathrm{Uniform}([0,\gamma/10])$ for $i\leq\lfloor d/2\rfloor$ (i.e. the $\mathcal{L}$ dimensions now have non-zero variance). Similarly, Figure 7 considers the same setting as Section 4.2 but with the first dimension of the input replaced with $y(\gamma+\xi)$ instead of $y\gamma$. As can be seen from the results, the introduction of a small amount of variance to these settings leaves model behavior effectively unchanged.
We also considered a similar low variance setup under the setting of Section 4.3. To do so, we modify the zero variance concatenated features with added mean-zero Gaussian noise in the training data. However, as we noted in the main paper, we found that unless the scaling of the added Gaussian noise was significantly less than $\gamma$ (i.e. $\gamma*10^{-3}$), the label smoothing and Mixup overfitting phenomenon was also exhibited by weight decay. This is possibly just a consequence of our experimental setup and/or the selected hyperparameters (i.e. perhaps this sensitivity is avoidable with some optimal choice of the weight decay hyperparameter), but we do not exhaustively explore this and instead leave a more comprehensive understanding of this phenomenon to follow-up work.

### v2

@@ 5 Conclusion
In this work, we have shown that label augmentation strategies such as label smoothing and Mixup exhibit a variance minimization effect that can lead to learning low variance, spurious correlations in the data even when there exists a better but higher variance feature. The main practical takeaway from this is that label smoothing and Mixup do not necessarily always improve the robustness of models, and their failure to do so may be indicative of certain identifying features in the training data (i.e. adversarially introduced patches or background pixels).
Limitations. The main theoretical limitation of work is that our results for general models are only loss lower bounds and do not characterize optimization dynamics. Our experiments that such dynamics for feedforward networks would still reflect our observations, however, and it would be an interesting follow-up to study these dynamics using the techniques of Zou et al. (2023) and Chidambaram et al. (2023). On the empirical side, our main limitation is that we have demonstrated our theory using only small-scale experiments; a comprehensive empirical evaluation of label smoothing and Mixup in larger scale spurious correlations settings guided by our theory would be a natural avenue for future work.
Broader Impacts. Although label augmentation are used to train and finetune large-scale models, our results concerning them in this work have mostly been theoretical. As a result, we do not anticipate any direct misuse of our results or any broader harmful impacts.
@@ Acknowledgements
Appendix A Omitted Proofs
A.1 Helper Lemma
The following lemma lower bounds the Jensen gap in terms of the variance of the random variable being considered. We will make repeated use of it in the proofs of the label smoothing and Mixup results.
Lemma A.1.
Let $\phi:\mathbb{R}^{d}\to\mathbb{R}$ be a twice-differentiable convex function satisfying $\gamma_{1}\mathrm{Id}\prec\nabla^{2}\phi\prec\gamma_{2}\mathrm{Id}$. Then for any square integrable random variable $X$ on $\mathbb{R}^{d}$ it follows that:
$\mathbb{E}[\phi(X)]-\phi(\mathbb{E}[X])\in\left[\frac{\gamma_{1}}{2}\mathrm{Var}(X),\frac{\gamma_{2}}{2}\mathrm{Var}(X)\right].$ (A.1)
@@ A.1 Helper Lemma
From the assumption of the lemma it follows that $\phi(x)-\gamma_{1}\norm{x}^{2}/2$ is convex and $\phi(x)-\gamma_{2}\norm{x}^{2}/2$ is concave. Applying Jensen’s inequality to each of these functions yields (A.1). ∎
A.2 Proofs for Section 3.1
See 3.3
Proof.
Let us decompose the unique minimizer $w^{*}$ as $w^{*}=\alpha_{1}u^{*}+\alpha_{2}v^{*}$, where $u^{*}$ and $v^{*}$ are orthonormal and satisfy $u^{*}_{\mathcal{H}}=0,v^{*}_{\mathcal{L}}=0$ (i.e. $u^{*}$ is a normalized version of the $\mathcal{L}$ components of $w^{*}$ and $v^{*}$ is the same but for the $\mathcal{H}$ components). We claim that $y\left\langle v^{*},x\right\rangle>y\left\langle u^{*},x\right\rangle$ for $\pi$-a.e. $(x,y)$. Indeed, if this were not the case then by the assumption on $\pi$ we could choose an orthonormal vector $z^{*}$ with $z^{*}_{\mathcal{L}}=0$ that satisfies $y\left\langle z^{*},x\right\rangle>y\left\langle u^{*},x\right\rangle$ for $\pi$-a.e. $(x,y)$ and decrease the loss of $w^{*}$ by replacing $v^{*}$ with $z^{*}$.
Now suppose that $\alpha_{1}>\alpha_{2}$. Then we claim that replacing $\alpha_{1}$ and $\alpha_{2}$ with $\gamma=\sqrt{(\alpha_{1}^{2}+\alpha_{2}^{2})/2}$ yields a solution with lower loss than $w^{*}$.
To see this, first observe that $2\gamma^{2}=\alpha_{1}^{2}+\alpha_{2}^{2}$, so that the norm of the modified solution is the same as $w^{*}$. This implies that the weight decay penalty term in $\ell_{\beta}$ is unchanged.
Furthermore, we have that $\gamma\in(\alpha_{2},\alpha_{1})$, and that $\gamma-\alpha_{2}>\alpha_{1}-\gamma$. This follows from the fact that $\gamma\geq(\alpha_{1}+\alpha_{2})/2$ by construction. This, combined with the fact that $y\left\langle v^{*},x\right\rangle>y\left\langle u^{*},x\right\rangle$, then implies for $\pi$-a.e. $(x,y)$ that:
$y\left\langle\gamma u^{*}+\gamma v^{*},x\right\rangle>y\left\langle\alpha_{1}u^{*}+\alpha_{2}v^{*},x\right\rangle.$ (A.2)
Which contradicts the optimality of $w^{*}$. Therefore, we must have $\alpha_{1}\leq\alpha_{2}$, which gives the desired result. ∎
See 3.4
@@
We follow the outline in the proof sketch. We first observe that there exists $M$ such that it suffices to consider $\norm{w}\leq M$; this is due to the fact that for $\alpha>0$, we have $\lim_{\norm{w}\to\infty}\absolutevalue{\ell_{\mathrm{LS},\alpha}(w)}=\infty$. With this in mind, let us use $Z$ to denote $w^{T}YX$ and use $\ell_{\mathrm{LS},\alpha}(Z)$ to denote the loss in terms of this quantity. Then it follows that
$\partialderivative{\ell_{\mathrm{LS},\alpha}(Z)}{Z}$ $=\mathbb{E}\left[\sigma(Z)-1+\alpha/2\right],$ (A.3)
[…]
where in both cases we applied the dominated convergence theorem, which is justified because $\ell_{\mathrm{LS},\alpha}$ is smooth in $Z$ with bounded derivatives. Now since $\norm{w}\leq M$ and the support of $X$ is compact, there exist $\gamma_{1}$ and $\gamma_{2}$ such that $2\partialderivative[2]{\ell_{\mathrm{LS},\alpha}(Z)}{Z}\in(\gamma_{1},\gamma_{2})$, which implies that $\ell_{\mathrm{LS},\alpha}$ is strongly convex in $Z$ and satisfies the conditions of Lemma A.1.
By Jensen’s inequality, we then have that:
$\ell_{\mathrm{LS},\alpha}(Z)\geq\left(\frac{\alpha}{2}-1\right)\log\sigma(\mathbb{E}[Z])-\frac{\alpha}{2}\log\sigma(-\mathbb{E}[Z]).$ (A.5)
Since $\mathbb{E}[YX_{\mathcal{L}}]\neq 0$ by Assumption 3.5, we can choose $w$ such that $w_{\mathcal{H}}=0$ and $\mathbb{E}[Z]=\sigma^{-1}(1-\alpha/2)$, which minimizes the RHS of (A.5). Let us use $\mathrm{OPT}$ to denote this minimum. Then by Lemma A.1, we have for any $w$ chosen as described:
$\ell_{\mathrm{LS},\alpha}(w)-\mathrm{OPT}\leq\frac{\gamma_{2}}{2}w^{T}\Sigma_{YX,\mathcal{L}}w.$ (A.6)
On the other hand, if $w^{\prime}$ is another solution satisfying $\norm{w^{\prime}_{\mathcal{H}}}>\epsilon$, then Lemma A.1 gives
$\ell_{\mathrm{LS},\alpha}(w^{\prime})-\mathrm{OPT}\geq\frac{\gamma_{1}\rho\epsilon^{2}}{2},$ (A.7)
from which it is clear that for appropriate $\epsilon$ we cannot have $w^{\prime}$ be a stationary point ($\rho$ above is the same as in Assumption 3.5). That we can take $\epsilon\to 0$ as $\norm{\Sigma_{YX,\mathcal{L}}}\to 0$ follows from (A.6). ∎
See 3.7
@@
Let us first outline the overall steps of the proof, and the differences with the label smoothing case.
  1. 1.
    We first show that the loss conditioned on $\lambda$ is strongly convex in $w^{\top}Z_{\lambda}$. The conditioning on $\lambda$ here is necessary because $\lambda$ is a random variable, unlike $\alpha$ in the label smoothing case. The overall goal here is to use the same argument as for label smoothing, i.e. show that we can achieve the optimal lower bound in terms of $\mathbb{E}[w^{\top}Z_{\lambda}]$ using only $w_{\mathcal{L}}$ and letting $\norm{\Sigma_{YX,\mathcal{L}}}\to 0$.
  2. 2.
    We cannot explicitly minimize the conditional loss like we did with label smoothing, since it is not possible with a fixed choice of $w$ to achieve $\sigma^{-1}(\mathbb{E}[w^{\top}Z_{\lambda}])=\lambda$ for every $\lambda$ simultaneously. Instead, we will show that a stationary point of the conditional loss exists that uses only the dimensions of $w_{\mathcal{L}}$.
  3. 3.
    Having shown the above, we can just reuse the same argument as before with Lemma A.1 to prove the desired result.
Let $\ell_{\mathrm{MIX},\lambda}$ denote $\ell_{\mathrm{MIX},\mathcal{D}_{\lambda}}$ with a fixed choice of $\lambda$ (i.e. after conditioning on $\lambda$), and let $R=w^{\top}Z_{\lambda}$. Then we can compute:
$\partialderivative{\ell_{\mathrm{MIX},\lambda}(R)}{R}$ $=\mathbb{E}\big{[}\lambda(\sigma(Y_{1}R)-1)Y_{1}+(1-\lambda)(\sigma(Y_{2}R)-1)Y_{2}\big{]},$ (A.8)
@@
where again we applied dominated convergence to the expectation with respect to $\pi$. Strong convexity follows from the same consideration as label smoothing; namely, we can consider $\norm{w}\leq M$ as $\ell_{\mathrm{MIX},\lambda}(R)\to\infty$ as $\norm{w}\to\infty$ so long as $\lambda$ is not 0 or 1 and $\pi_{Y}(y)>0$, and this consequently implies (A.9) is lower bounded by some positive real number.
Now by conditional Jensen’s inequality, we obtain the following lower bound for $\ell_{\mathrm{MIX},\mathcal{D}_{\lambda}}$:
$\ell_{\mathrm{MIX},\mathcal{D}_{\lambda}}(w)\geq\mathbb{E}_{\lambda}\left[\mathbb{E}_{Y_{1},Y_{2}}\left[\ell_{\mathrm{MIX},\lambda}(\mathbb{E}\left[R\mid\lambda,Y_{1},Y_{2}\right])\right]\right].$ (A.10)
We now show that it is possible to minimize this lower bound while taking $w_{\mathcal{H}}=0$. This is more difficult than it was in the label smoothing case, because it is no longer obvious that $\mathbb{E}[YX_{\mathcal{L}}]\neq 0$ is sufficient for minimizing the RHS of (A.10) due to the expectation with respect to $\lambda$. However, we can show the existence of a stationary point with $w_{\mathcal{H}}=0$, even though we cannot provide an explicit construction.
The idea is to consider the limiting behavior of (A.10) as we take the values of $w_{\mathcal{L}}$ to $\infty$- ∞ and $\infty$. Note that we can take this limit into the expectation with respect to $\lambda$ by dominated convergence again. Let us consider the gradient with respect to $w$ of the RHS of (A.10). To make notation manageable, we will use $a_{1}=\mathbb{E}[X\mid Y=1],a_{2}=\mathbb{E}[X\mid Y=-1]$, and $a_{3}=\mathbb{E}[Z_{\lambda}\mid\lambda,Y_{1}=1,Y_{2}=-1]$. We can then explicitly write out the gradient as the expectation with respect to $\lambda$ of the sum of the following three terms (considering the different cases for $Y_{1},Y_{2}$):
$\gradient_{w}\mathbb{E}_{Y_{1},Y_{2}}\left[\ell_{\mathrm{MIX},\lambda}(\mathbb{E}\left[R\mid\lambda,Y_{1},Y_{2}\right])\right]$ $=\pi_{Y}(1)^{2}\left(\sigma(w^{\top}a_{1})-1\right)a_{1}$
$\quad-\pi_{Y}(-1)^{2}\left(\sigma(-w^{\top}a_{2})-1\right)a_{2}$
$\quad+2\pi_{Y}(1)\pi_{Y}(-1)\bigg{(}\lambda\big{(}\sigma(w^{\top}a_{3})-1\big{)}$
$\quad\quad\quad-(1-\lambda)\big{(}\sigma(-w^{\top}a_{3})-1\big{)}\bigg{)}a_{3}.$ (A.11)
The first two lines above are obtained from the fact that we can combine terms when $Y_{1}=Y_{2}$, and the last line is by symmetry. Now we recall that by assumption $\mathbb{E}[YX_{\mathcal{L}}]\neq 0$ and $\mathbb{E}[X]=0$. Thus, WLOG, we can assume that $\mathbb{E}[X_{\mathcal{L}}\mid Y=1]_{i}>0$ and $\mathbb{E}[X_{\mathcal{L}}\mid Y=-1]_{i}<0$ for each index $i$.
With this in mind, we consider first the case of what happens when the entries $w_{\mathcal{L}}\to\infty$. Since $w^{\top}a_{1}>0$ and $w^{\top}a_{2}<0$, the first two terms in (A.11) vanish (independent of $\lambda$). On the other hand, for the third term, there are two cases to consider. Depending on $\lambda$, we have that the entries of $a_{3}$ are either strictly negative or strictly positive, with the exceptional case of $\lambda=\pi_{Y}(1)$ in which $a_{3}=0$. If the entries of $a_{3}$ are strictly negative, then $\big{(}\sigma(-w^{\top}a_{3})-1\big{)}\to 0$ and the coefficient becomes negative, so the third term is positive. Similarly, if $a_{3}$ is strictly positive, the coefficient is positive and the third term is still positive. Thus, as $w_{\mathcal{L}}\to\infty$ every entry of (A.11) is positive.
@@
Now we have basically arrived at the same stage as the end of the label smoothing proof. By taking $\norm{\Sigma_{YX,\mathcal{L}}}\to 0$ we can get arbitrary concentration around this optimal $R$, and by the same logic as the label smoothing proof the result follows. ∎
A.3 Proofs for Section 3.2
See 3.9
Proof.
Let us first define:
$\ell_{\mathrm{LS},\alpha,y}(g)=-\mathbb{E}\bigg{[}(1-\alpha)\log g^{y}(X)+\frac{\alpha}{k}\sum_{i=1}^{k}\log g^{i}(X)\mid Y=y\bigg{]}.$ (A.12)
Then we can decompose $\ell_{\mathrm{LS},\alpha}(g)$ as follows:
$\ell_{\mathrm{LS},\alpha}(g)=\sum_{y=1}^{k}\pi_{Y}(y)\ell_{\mathrm{LS},\alpha,y}(g).$ (A.13)
Once again, since we can restrict our attention to $g(X)\in[\gamma,1-\gamma]$ for some $\gamma$ (as the loss goes to infinity if $g(X)=1$ on a set of positive $\pi_{X}$-measure with $\alpha>0$), it is easy to verify that (A.12) is strongly convex in $g(X)$. The desired result then follows by applying Lemma A.1 with the regular conditional distribution $g(X)\mid Y=y$ for each term in (A.13). ∎
See 3.10
Proof.
The proof follows an identical structure to that of Proposition 3.9. In particular, we again define the following conditional loss:
$\ell_{\mathrm{MIX},\lambda,y_{1},y_{2}}=-\mathbb{E}[\lambda\log g^{y_{1}}(Z_{\lambda})+(1-\lambda)\log g^{y_{2}}(Z_{\lambda})\mid y_{1},y_{2},\lambda].$ (A.14)
And we can then decompose $\ell_{\mathrm{MIX},\mathcal{D}_{\lambda}}$ as:
$\ell_{\mathrm{MIX},\lambda}(g)=\sum_{y_{1}=1}^{k}\sum_{y_{2}=1}^{k}\pi_{Y}(y_{1})\pi_{Y}(y_{2})\mathbb{E}_{\lambda\sim\mathcal{D}_{\lambda}}[\ell_{\mathrm{MIX},\lambda,y_{1},y_{2}}(g)].$ (A.15)
Since $\mathcal{D}_{\lambda}$ is not a point mass on 00 or $1$, we can restrict ourselves to $g(X)\in[\gamma,1-\gamma]$ for some $\gamma$ as before, and strong convexity in $g(X)$ of the conditional loss (A.14) again follows. We then apply Lemma A.1 to obtain the result. ∎